# Random regular graph states are complex at almost any depth

### v1

@@ 5.3 Structured simulation algorithms
The upper bound on the geometric entanglement derived above is strong enough to break the simulation algorithm derived in [24]. However, better simulation algorithms might be found by exploiting the extra structure that graph states provide. In particular, we can think of any MBQC procedure as a sequence of measurements in the eigenbases of the $X,Y,X-Y,X+Y$ operators [70]. On a graph state the first three measurements can be simulated classically in polynomial time. We can thus envision an improved simulation algorithm where we classically simulate the $X,Y,Z$ measurements and simulate the $X+Y,X-Y$ measurements by coin flips. The efficacy of this algorithm depends critically on how the overlap with a random graph state fluctuates with respect to tensor products of the eigenstates of $X+Y,X-Y$. As these states are highly magical, it is possible that their overlap with stabilizer states fluctuates much less than the maximal overlap with arbitrary stabilizer states (which as we saw above, is dominated by contributions from product stabilizer states). We believe that this is not the case (and thus that this algorithm does not work), but can not prove it as of yet. We think this is an interesting question in its own right, so we leave it as a conjecture.
@@
Labeling the set of eigenstates of $X+Y,X-Y$ as $\mathcal{T}$, there exist $c,C>0$ s.t.
$\mathbb{P}\left[\max_{\ket{\alpha}\in\mathcal{T}^{\otimes n}}|\langle\alpha\ket{G}|^{2}\geq 2^{-n+n^{c}}\right]\geq C,$ (238)

### v2

@@ 5.3 Structured simulation algorithms
The upper bound on the geometric entanglement derived above is strong enough to break the simulation algorithm derived in [38]. However, better simulation algorithms might be found by exploiting the extra structure that graph states provide. In particular, we can think of any MBQC procedure as a sequence of measurements in the eigenbases of the X,Y,XY,X+YX,Y,X-Y,X+Y operators [90]. On a graph state the first three measurements can be simulated classically in polynomial time. We can thus envision an improved simulation algorithm where we classically simulate the $X,Y,Z$ measurements and simulate the X+Y,XYX+Y,X-Y measurements by coin flips. The efficacy of this algorithm depends critically on how the overlap with a random graph state fluctuates with respect to tensor products of the eigenstates of X+Y,XYX+Y,X-Y. As these states are highly magical, it is possible that their overlap with stabilizer states fluctuates much less than the maximal overlap with arbitrary stabilizer states (which as we saw above, is dominated by contributions from product stabilizer states). We believe that this is not the case (and thus that this algorithm does not work), but can not prove it as of yet. We think this is an interesting question in its own right, so we leave it as a conjecture.
@@
Labeling the set of eigenstates of X+Y,XYX+Y,X-Y as $\mathcal{T}$, there exist $c,C>0$ s.t.
[…]